### v2

@@ 4 Next-token prediction capacity
From here on out, we will consider the finite set in our probabilistic language space to be $[\omega]=\{1,\ldots,\omega\}$ for some $\omega\in\mathbb{N}$. We call $[\omega]$ a vocabulary (it can also be called an alphabet) and we call its elements tokens. The tokens may represent letters, words, or something in between, as in Byte Pair Encoding (Gage, 1994). We assume that there is an underlying probabilistic language space of the first kind, $([\omega],A,p)$, but that we are only given the values of $p$ on $[\omega]\times\{\alpha^{1},\ldots,\alpha^{n}\}\subset[\omega]\times A$ for some $n\in\mathbb{N}$. We call each $\alpha^{i}$ a context and we call $[p(\gamma\mid\alpha)]_{\gamma\in[\omega]}\in\Delta^{\omega-1}$ its corresponding next-token distribution. The task of next-token prediction is for a model of the form $\cup_{t=0}^{\infty}[\omega]^{t}\to\mathrm{ri}\Delta^{\omega-1}$ to correctly predict next-token distributions given contexts (this implies that if any of the next-token distributions come from the relative boundary of $\Delta^{\omega-1}$, then the model can still approximate them up to arbitrary precision). In particular, the largest $n$ such that it can do so exactly is what we call the next-token prediction capacity. We will formally define this, then prove an upper bound.
Definition 4.1.
Let $k,\omega\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\omega]^{t}\to\mathrm{ri}\Delta^{\omega-1}.$
Then the next-token prediction capacity of the hypothesis class $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is the largest $n\in\mathbb{N}$ such that for every distinct list $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=0}^{\infty}[\omega]^{t}$ and every list $y_{1},\ldots,y_{n}\in\mathrm{ri}\Delta^{\omega-1}$ there exists $\theta\in\mathbb{R}^{k}$ such that $y_{i}=h_{\theta}(\alpha^{i})\leavevmode\nobreak\ \forall i\in[n]$.
Lemma 4.2.
Let $k,\omega\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\omega]^{t}\to\mathrm{ri}\Delta^{\omega-1}.$
Assume that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable for all $\alpha\in\cup_{t=0}^{\infty}[\omega]^{t}$. Then the next-token prediction capacity of $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is less than or equal to $k/(\omega-1)$.
@@
Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\omega]^{t}$ be distinct. Define $F:\mathbb{R}^{k}\to(\mathrm{ri}\Delta^{\omega-1})^{n}:\theta\mapsto[h_{\theta}(\alpha^{i})]_{i\in[n]}$. $F$ is continuously differentiable, so we can apply Sard’s theorem (Sard, 1942). Note that $\mathrm{ri}\Delta^{\omega-1}$ is an $(\omega-1)$-dimensional manifold. Thus, if $k<n(\omega-1)$, then the image of $F$ has measure zero, hence $F$ is not surjective, proving the theorem. ∎
In practice, we are generally given a corpus of $N$ document sequences $\beta^{1},\ldots,\beta^{N}\in\cup_{t=1}^{\infty}[\omega]^{t}$. We can think of these as finite-time samples of the underlying stochastic process. Note that each $\beta^{j}$ has $|\beta^{j}|$ contexts $(),(\beta_{1}^{j}),(\beta_{1}^{j},\beta_{2}^{j}),\ldots$. But, among the corpus, there may be repeated contexts, so we enumerate $\{(\beta_{1}^{j},\ldots,\beta_{t}^{j})\mid j\in[N],t<|\beta^{j}|\}$ as $\{\alpha^{1},\ldots,\alpha^{n}\}$. We call $\alpha^{1},\ldots,\alpha^{n}$ the unique contexts of the corpus. For all $\alpha\in\cup_{t=0}^{\infty}[\omega]^{t}$, define $c(\alpha)=|\{j\in[N]\mid\alpha=(\beta_{1}^{j},\ldots,\beta_{|\alpha|}^{j})\}|$. This is the number of times that the context $\alpha$ appears in the corpus. In particular, $c(\leavevmode\nobreak\ )=N$. Now we are ready to define the underlying $p\in\mathbb{L}_{1}$ (where $\mathbb{L}_{1}$ is as defined in Proposition 3.4) which we denote by $\hat{p}$ to indicate that it comes from a corpus:
$\hat{p}(\gamma\mid\alpha)=c(\alpha,\gamma)/c(\alpha)\leavevmode\nobreak\ \forall\alpha\in\cup_{t=0}^{\infty}[\omega]^{t}\text{ s.t. }c(\alpha)\neq 0,\leavevmode\nobreak\ \forall\gamma\in[\omega].$ (1)
By the usual abuse of notation, let $\hat{p}$ denote $\Phi_{12}(\hat{p})$ (where $\Phi_{12}$ is as defined in Proposition 3.4). Then $\hat{p}$ satisfies $\hat{p}(\alpha)=c(\alpha)/N\leavevmode\nobreak\ \forall\alpha\in\cup_{t=1}^{\infty}[\omega]^{t}$.
@@
To see how it relates to the usual notion of the cross-entropy of a probability mass function $q:\mathcal{X}\to[0,1]$ relative to another probability mass function $p:\mathcal{X}\to[0,1]$, $H(p,q)\coloneqq\sum_{x\in\mathcal{X}}p(x)\log q(x)$, let $T$ denote $\max_{j\in[N]}|\beta^{j}|$ and observe
$\mathcal{L}(\beta^{1},\ldots,\beta^{N},q)$ $=-\sum_{t=1}^{T}\sum_{\alpha\in[\omega]^{t-1}}\sum_{\gamma=1}^{\omega}c(\alpha,\gamma)\log q(\gamma\mid\alpha)=-N\sum_{t=1}^{T}\sum_{\alpha\in[\omega]^{t-1}}\sum_{\gamma=1}^{\omega}\hat{p}(\alpha,\gamma)\log q(\gamma\mid\alpha)$ (3)
$=-\sum_{t=1}^{T}\sum_{\alpha\in[\omega]^{t-1}}c(\alpha)\sum_{\gamma=1}^{\omega}\hat{p}(\gamma\mid\alpha)\log q(\gamma\mid\alpha)=-\sum_{t=1}^{T}\sum_{\alpha\in[\omega]^{t-1}}c(\alpha)H(\hat{p}(\cdot\mid\alpha),q(\cdot\mid\alpha)).$ (4)
Eq. (3) shows that $\mathcal{L}$ is the sum of the $t$-gram entropies, as defined in Shannon (1948), up to $t=T-1$. Eq. (4) shows that $\mathcal{L}$ is a positively weighted sum of entropies. Thus, by Gibb’s inequality,
$\mathcal{L}(\beta^{1},\ldots,\beta^{N},\hat{p})\leq\mathcal{L}(\beta^{1},\ldots,\beta^{N},q)\text{ with equality iff }\hat{p}=q.$ (5)
This makes the connection between optimization and interpolation sufficiently clear to make the following definition, from which follows a lower bound.
Definition 4.3.
Let $k,\omega\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\omega]^{t}\to\mathrm{ri}\Delta^{\omega-1}.$
Then the empirical next-token prediction capacity of the hypothesis class $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is the largest $n\in\mathbb{N}$ such that for every $N\in\mathbb{N}$ and every list $\beta^{1},\ldots,\beta^{N}\in\cup_{t=1}^{\infty}[\omega]^{t}$ such that |$|\{(\beta_{1}^{j},\ldots,\beta_{t}^{j})\mid j\in[N],t<|\beta^{j}|\}|\leq n$ the following holds:
@@
Lemma 4.4.
Let $k,\omega\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\omega]^{t}\to\mathrm{ri}\Delta^{\omega-1}.$
@@
Thus, any lower bound on the next-token prediction capacity of a model is a lower bound on the empirical next-token prediction capacity of the model as well. On the other hand, we do not yet have a general upper bound on the empirical next-token prediction capacity, so we prove one in the following theorem. As $N$ increases, we show that $[[\hat{p}(\gamma\mid\alpha^{i})]_{\gamma\in[\omega]}]_{i\in[n]}$ can approximate any $Y\in(\Delta^{\omega-1})^{n}$. But, this is not enough to apply Lemma 4.2 since the image of $F$ in the proof, despite being measure zero, may still be dense. To guarantee that it is not dense, we have to further assume that $\theta\mapsto h_{\theta}(\alpha)$ is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=0}^{\infty}[\omega]^{t}$. As an example, this is satisfied by the transformer model described in the next section for activations such as tanh, arctan, and GELU.
Theorem 4.5.
Let $k,\omega\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\omega]^{t}\to\mathrm{ri}\Delta^{\omega-1}.$
Assume that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=0}^{\infty}[\omega]^{t}$. Then the empirical next-token prediction capacity of $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is less than or equal to $(2+1/(\omega-1))k/(\omega-1)+2$.
@@
Set $n^{\prime}=\lfloor k/(\omega-1)+1\rfloor$ and $T=\lceil\log(n^{\prime})/\log(\omega)\rceil$. Define $\hat{p}$ and $\mathcal{L}$ as in Eqs. (1) and (2) respectively. We will show that there exists $N\geq n^{\prime}$ and $\beta^{1},\ldots,\beta^{N}\in[\omega]^{N+1}$ such that the number of unique contexts is less than or equal to $(2+1/(\omega-1))k/(\omega-1)+2$ and
@@
First, the size of $[\omega]^{T}$ is $\omega^{T}\geq n^{\prime}\geq\omega^{T-1}$, so there exists a distinct list $\alpha^{1},\ldots,\alpha^{n^{\prime}}\in[\omega]^{T}$. Choose such an $\alpha^{1},\ldots,\alpha^{n^{\prime}}$. Let $\beta^{1},\ldots,\beta^{n^{\prime}}\in[\omega]^{T+1}$ such that $\beta^{i}[1:T]=\alpha^{i}\leavevmode\nobreak\ \forall i\in[n^{\prime}]$. Then we can bound the total number of unique contexts
$n$ $\coloneqq n^{\prime}+\sum_{t=0}^{T-1}\omega^{t}=n^{\prime}+\frac{\omega^{T}-1}{\omega-1}\leq n^{\prime}+\frac{\omega n^{\prime}-1}{\omega-1}=\frac{2\omega-1}{\omega-1}n^{\prime}-\frac{1}{\omega-1}\leq\frac{2\omega-1}{\omega-1}\left(\frac{k}{\omega-1}+1\right)-\frac{1}{\omega-1}.$
@@
Define $F:\mathbb{R}^{k}\to(\mathrm{ri}\Delta^{\omega-1})^{n^{\prime}}:\theta\mapsto[h_{\theta}(\alpha^{i})]_{i\in[n^{\prime}]}$. $F$ is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable. Since $k<n^{\prime}(\omega-1)$, the image of $F$ is measure zero by Sard’s theorem, hence has empty interior. So, by Corollary B.4 of Sontag (1996), it is nowhere dense, hence not dense. Let $Y\in(\Delta^{\omega-1})^{n^{\prime}}\backslash\overline{F(\mathbb{R}^{k})}$. Note that we can include more sequences $\beta^{n^{\prime}+1},\ldots,\beta^{N}$ without adding new contexts. Furthermore, we have control over the last token in every sequence, so by increasing $N$, we can make $[[\hat{p}(\gamma\mid\alpha^{i})]_{\gamma\in[\omega]}]_{i\in[n^{\prime}]}$ approximate $Y$ to arbitrary precision. Thus, since $Y\notin\overline{F(\mathbb{R}^{k})}$, there exists $N\geq n^{\prime}$ and $\beta^{1},\ldots,\beta^{N}\in[\omega]^{T+1}$ for which $[[\hat{p}(\gamma\mid\alpha^{i})]_{\gamma\in[\omega]}]_{i\in[n^{\prime}]}\notin\overline{F(\mathbb{R}^{k})}$. So, the conclusion follows from Eq. (5). ∎
@@
In this section, we have defined next-token prediction capacity and empirical next-token prediction capacity and upper bounded them for general models. The rest of the paper will focus on lower bounding the next-token prediction capacity of the one-layer multi-head decoder-only transformer model, which we define in the next section. Before we move on though, we end the section with two remarks on inexact interpolation and finite precision.
Remark 4.6.
@@
Both Lemma 4.2 and Theorem 4.5 apply even for inexact interpolation. While Lemma 4.2 only uses that $F$, as defined in the proof, is not surjective, the proof actually shows that the image of $F$ has measure zero. If $\theta\mapsto h_{\theta}(\alpha)$ is additionally assumed to be $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=0}^{\infty}[\omega]^{t}$, as it is in Theorem 4.5, then the image of $F$ is nowhere dense. Similarly, the image of $F$ defined in the proof of Theorem 4.5 is shown to be nowhere dense. So, for each choice of $k$ and $\omega$, there exists $\epsilon>0$ such that the “$\epsilon$-approximate” next-token prediction capacity and “$\epsilon$-approximate” empirical next-token prediction capacity are both upper bounded by $O(k/\omega)$.
Remark 4.7.
@@
Any next-token prediction capacity lower bound for a hypothesis class will apply even if the model is constrained to finite precision. To see this, suppose we have a hypothesis class $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ such that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable for all $\alpha\in\cup_{t=0}^{\infty}[\omega]^{t}$. If we assume a maximum context length $T$ and restrict $\theta$ to a compact set, then the $\theta\mapsto h_{\theta}(\alpha)$ are uniformly Lipschitz continuous over $\alpha\in\cup_{t=0}^{\infty}[\omega]^{t}$, so there exists $\delta>0$ such that any next-token prediction capacity lower bound also holds as an “$\epsilon$-approximate” next-token prediction capacity lower bound for the hypothesis class $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ restricted to $\delta$-precision.

### v3

@@ 4 Next-token Prediction Capacity
From here on out, we will consider the finite set in our probabilistic language space to be $[\zeta]=\{1,\ldots,\zeta\}$ for some $\zeta\in\mathbb{N}$. We call $[\zeta]$ a vocabulary (it can also be called an alphabet) and we call its elements tokens. The tokens may represent letters, words, or something in between, as in Byte Pair Encoding [54]. We assume that there is an underlying probabilistic language space of the first kind, $([\zeta],A,p)$, but that we are only given the values of $p$ on $[\zeta]\times\{\alpha^{1},\ldots,\alpha^{n}\}\subset[\zeta]\times A$ for some $n\in\mathbb{N}$. We call each $\alpha^{i}$ a context and we call $[p(\gamma\mid\alpha)]_{\gamma\in[\zeta]}\in\Delta^{\zeta-1}$ its corresponding next-token distribution. The task of next-token prediction is for a model of the form $\cup_{t=0}^{\infty}[\zeta]^{t}\to\mathrm{ri}\Delta^{\zeta-1}$ to correctly predict next-token distributions given contexts (this implies that if any of the next-token distributions come from the relative boundary of $\Delta^{\zeta-1}$, then the model can still approximate them up to arbitrary precision). In particular, the largest $n$ such that it can do so exactly is what we call the next-token prediction capacity. We will formally define this, then prove an upper bound.
Definition 7.
Let $k,\zeta\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\zeta]^{t}\to\mathrm{ri}\Delta^{\zeta-1}.$
Then the next-token prediction capacity of the hypothesis class $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is the largest $n\in\mathbb{N}$ such that for every distinct list $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=0}^{\infty}[\zeta]^{t}$ and every list $y_{1},\ldots,y_{n}\in\mathrm{ri}\Delta^{\zeta-1}$ there exists $\theta\in\mathbb{R}^{k}$ such that $y_{i}=h_{\theta}(\alpha^{i})\leavevmode\nobreak\ \forall i\in[n]$.
Example 8.
Let 1 represent ‘I,’ 2 represent ‘love,’ 3 represent ‘hate,’ 4 represent ‘cats,’ and 5 represent ‘dogs.’ Let $\alpha^{1}=(\leavevmode\nobreak\ )$, $\alpha^{2}=(1)$, $\alpha^{3}=(1,2)$, and $\alpha^{4}=(1,3)$. Let $y_{1}=(1,0,0,0,0)$, $y_{2}=(0,2/3,1/3,0,0)$, $y_{3}=(0,0,0,1/2,1/2)$, and $y_{4}=(0,0,0,1,0)$. This fully defines the first three terms of a stochastic sequence. The first term is ‘I’ with probability 1. The second term is ‘love’ with probability 2/3, ‘hate’ otherwise. If the first two terms are ‘I’ and ‘love,’ then the third term is ‘cats’ with probability 1/2, ‘dogs’ otherwise. If the first two terms are ‘I’ and ‘hate,’ then the third term is ‘cats’ with probability 1.
Lemma 9.
Let $k,\zeta\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\zeta]^{t}\to\mathrm{ri}\Delta^{\zeta-1}.$
Assume that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable for all $\alpha\in\cup_{t=0}^{\infty}[\zeta]^{t}$. Then the next-token prediction capacity of $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is less than or equal to $k/(\zeta-1)$.
@@
Let $\alpha^{1},\ldots,\alpha^{n}\in\cup_{t=1}^{\infty}[\zeta]^{t}$ be distinct. Define $F:\mathbb{R}^{k}\to(\mathrm{ri}\Delta^{\zeta-1})^{n}:\theta\mapsto[h_{\theta}(\alpha^{i})]_{i\in[n]}$. $F$ is continuously differentiable, so we can apply Sard’s theorem (Lemma 1). Note that $\mathrm{ri}\Delta^{\zeta-1}$ is an $(\zeta-1)$-dimensional manifold. Thus, if $k<n(\zeta-1)$, then the image of $F$ has measure zero in $(\mathrm{ri}\Delta^{\zeta-1})^{n}$, hence $F$ is not surjective, proving the theorem. ∎
In practice, we are generally given a corpus of $N$ document sequences $\beta^{1},\ldots,\beta^{N}\in\cup_{t=1}^{\infty}[\zeta]^{t}$. We can think of these as finite-time samples of the underlying stochastic process. Note that each $\beta^{j}$ has $|\beta^{j}|$ contexts $(),(\beta_{1}^{j}),(\beta_{1}^{j},\beta_{2}^{j}),\ldots$. But, among the corpus, there may be repeated contexts, so we enumerate $\{(\beta_{1}^{j},\ldots,\beta_{t}^{j})\mid j\in[N],t<|\beta^{j}|\}$ as $\{\alpha^{1},\ldots,\alpha^{n}\}$. We call $\alpha^{1},\ldots,\alpha^{n}$ the unique contexts of the corpus. For all $\alpha\in\cup_{t=0}^{\infty}[\zeta]^{t}$, define $c(\alpha)=|\{j\in[N]\mid\alpha=(\beta_{1}^{j},\ldots,\beta_{|\alpha|}^{j})\}|$. This is the number of times that the context $\alpha$ appears in the corpus. In particular, $c(\leavevmode\nobreak\ )=N$. Now we are ready to define the underlying $p\in\mathbb{L}_{1}$ (where $\mathbb{L}_{1}$ is as defined in Proposition 5) which we denote by $\hat{p}$ to indicate that it comes from a corpus: for all $\gamma\in[\zeta]$, for all $\forall\alpha\in\cup_{t=0}^{\infty}[\zeta]^{t}\text{ s.t. }c(\alpha)\neq 0$,
$\hat{p}(\gamma\mid\alpha)=c(\alpha,\gamma)/c(\alpha).$ (1)
By the usual abuse of notation, let $\hat{p}$ denote $\Phi_{12}(\hat{p})$ (where $\Phi_{12}$ is as defined in Proposition 5). Then $\hat{p}$ satisfies $\hat{p}(\alpha)=c(\alpha)/N\leavevmode\nobreak\ \forall\alpha\in\cup_{t=1}^{\infty}[\zeta]^{t}$.
@@
To see how it relates to the usual notion of the cross-entropy of a probability mass function $q:\mathcal{X}\to[0,1]$ relative to another probability mass function $p:\mathcal{X}\to[0,1]$, $H(p,q)\coloneqq-\sum_{x\in\mathcal{X}}p(x)\log q(x)$, let $T$ denote $\max_{j\in[N]}|\beta^{j}|$ and observe
$,\displaystyle\mathcal{L}(\beta^{1},$$\ldots,\beta^{N},q)$
$=-\sum_{t=1}^{T}\sum_{\alpha\in[\zeta]^{t-1}}\sum_{\gamma=1}^{\zeta}c(\alpha,\gamma)\log q(\gamma\mid\alpha)$
$=-N\sum_{t=1}^{T}\sum_{\alpha\in[\zeta]^{t-1}}\sum_{\gamma=1}^{\zeta}\hat{p}(\alpha,\gamma)\log q(\gamma\mid\alpha)$
$=-\sum_{t=1}^{T}\sum_{\alpha\in[\zeta]^{t-1}}c(\alpha)\sum_{\gamma=1}^{\zeta}\hat{p}(\gamma\mid\alpha)\log q(\gamma\mid\alpha)$
$=\sum_{t=1}^{T}\sum_{\alpha\in[\zeta]^{t-1}}c(\alpha)H(\hat{p}(\cdot\mid\alpha),q(\cdot\mid\alpha)).$
The second equality shows that $\mathcal{L}$ is the sum of the $t$-gram entropies, as defined in Shannon [6], up to $t=T-1$. The last equality shows that $\mathcal{L}$ is a positively weighted sum of entropies. Thus, by Gibb’s inequality,
$\mathcal{L}(\beta^{1},\ldots,\beta^{N},\hat{p})\leq\mathcal{L}(\beta^{1},\ldots,\beta^{N},q)$ (3)
with equality if and only if $\hat{p}=q$. This makes the connection between optimization and interpolation sufficiently clear to make the following definition, from which follows a lower bound.
Definition 10.
Let $k,\zeta\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\zeta]^{t}\to\mathrm{ri}\Delta^{\zeta-1}.$
Then the empirical next-token prediction capacity of the hypothesis class $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is the largest $n\in\mathbb{N}$ such that for every $N\in\mathbb{N}$ and every list $\beta^{1},\ldots,\beta^{N}\in\cup_{t=1}^{\infty}[\zeta]^{t}$ such that |$|\{(\beta_{1}^{j},\ldots,\beta_{t}^{j})\mid j\in[N],t<|\beta^{j}|\}|\leq n$ the following holds:
@@
Example 11.
Consider a text corpus made up of three documents, ‘I love cats,’ ‘I love dogs,’ and ‘I hate cats.’ If we ignore spaces and tokenize as in Example 8, then we get $\beta^{1}=(1,2,4)$, $\beta^{2}=(1,2,5)$, and $\beta^{3}=(1,3,4)$. Then $\hat{p}(\cdot\mid\alpha^{i})=y_{i}$ for each $i\in[4]$, where the $\alpha^{i}$ and $y_{i}$ are as defined in Example 8. In this case, $\mathcal{L}(\beta^{1},\beta^{2},\beta^{3},\hat{p})=3\log(3)$.
Lemma 12.
Let $k,\zeta\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\zeta]^{t}\to\mathrm{ri}\Delta^{\zeta-1}.$
@@
Thus, any lower bound on the next-token prediction capacity of a model is a lower bound on the empirical next-token prediction capacity of the model as well. On the other hand, we do not yet have a general upper bound on the empirical next-token prediction capacity, so we prove one in the following theorem. As $N$ increases, we show that $[[\hat{p}(\gamma\mid\alpha^{i})]_{\gamma\in[\zeta]}]_{i\in[n]}$ can approximate any $Y\in(\Delta^{\zeta-1})^{n}$. But, this is not enough to apply Lemma 9 since the image of $F$ in the proof, despite being measure zero, may still be dense. To guarantee that it is not dense, we have to further assume that $\theta\mapsto h_{\theta}(\alpha)$ is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=0}^{\infty}[\zeta]^{t}$. As explained in Section 2, $\phi$, tanh, arctan, erf, and GELU are $(\mathbb{R}_{\mathrm{an}},\exp)$-definable. Furthermore, if $h_{\theta}$ is the transformer model defined in the next section, and if the activation function is a $(\mathbb{R}_{\mathrm{an}},\exp)$-definable function, such as tanh, arctan, or GELU, then $h_{\theta}$ is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable.
Theorem 13.
Let $k,\zeta\in\mathbb{N}$. For all $\theta\in\mathbb{R}^{k}$, let
$h_{\theta}:\cup_{t=0}^{\infty}[\zeta]^{t}\to\mathrm{ri}\Delta^{\zeta-1}.$
Assume that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=0}^{\infty}[\zeta]^{t}$. Then the empirical next-token prediction capacity of $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ is less than or equal to $(2+1/(\zeta-1))k/(\zeta-1)+2$.
@@
Set $n^{\prime}=\lfloor k/(\zeta-1)+1\rfloor$ and $T=\lceil\log(n^{\prime})/\log(\zeta)\rceil$. Define $\hat{p}$ and $\mathcal{L}$ as in Eqs. (1) and (2) respectively. We will show that there exists $N\geq n^{\prime}$ and $\beta^{1},\ldots,\beta^{N}\in[\zeta]^{N+1}$ such that the number of unique contexts is less than or equal to $(2+1/(\zeta-1))k/(\zeta-1)+2$ and
@@
First, the size of $[\zeta]^{T}$ is $\zeta^{T}\geq n^{\prime}\geq\zeta^{T-1}$, so there exists a distinct list $\alpha^{1},\ldots,\alpha^{n^{\prime}}\in[\zeta]^{T}$. Choose such an $\alpha^{1},\ldots,\alpha^{n^{\prime}}$. Let $\beta^{1},\ldots,\beta^{n^{\prime}}\in[\zeta]^{T+1}$ such that $\beta^{i}[1:T]=\alpha^{i}\leavevmode\nobreak\ \forall i\in[n^{\prime}]$. Then we can bound the total number of unique contexts
$n$ $\coloneqq n^{\prime}+\sum_{t=0}^{T-1}\zeta^{t}$
$=n^{\prime}+\frac{\zeta^{T}-1}{\zeta-1}$
$\leq n^{\prime}+\frac{\zeta n^{\prime}-1}{\zeta-1}$
$=\frac{2\zeta-1}{\zeta-1}n^{\prime}-\frac{1}{\zeta-1}$
$\leq\frac{2\zeta-1}{\zeta-1}\left(\frac{k}{\zeta-1}+1\right)-\frac{1}{\zeta-1}.$
@@
Define $F:\mathbb{R}^{k}\to(\mathrm{ri}\Delta^{\zeta-1})^{n^{\prime}}:\theta\mapsto[h_{\theta}(\alpha^{i})]_{i\in[n^{\prime}]}$. $F$ is continuously differentiable and $(\mathbb{R}_{\mathrm{an}},\exp)$-definable. Since $k<n^{\prime}(\zeta-1)$, the image of $F$ is measure zero in $(\mathrm{ri}\Delta^{\zeta-1})^{n^{\prime}}$ by Sard’s theorem (Lemma 1), hence has empty interior. So, by Corollary B.4 of Sontag [48], it is nowhere dense, hence not dense. Let $Y\in(\Delta^{\zeta-1})^{n^{\prime}}\backslash\overline{F(\mathbb{R}^{k})}$. Note that we can include more sequences $\beta^{n^{\prime}+1},\ldots,\beta^{N}$ without adding new contexts. Furthermore, we have control over the last token in every sequence, so by increasing $N$, we can make $[[\hat{p}(\gamma\mid\alpha^{i})]_{\gamma\in[\zeta]}]_{i\in[n^{\prime}]}$ approximate $Y$ to arbitrary precision. Thus, since $Y\notin\overline{F(\mathbb{R}^{k})}$, there exists $N\geq n^{\prime}$ and $\beta^{1},\ldots,\beta^{N}\in[\zeta]^{T+1}$ for which $[[\hat{p}(\gamma\mid\alpha^{i})]_{\gamma\in[\zeta]}]_{i\in[n^{\prime}]}\notin\overline{F(\mathbb{R}^{k})}$. So, the conclusion follows from Eq. (3). ∎
@@
In this section, we have defined next-token prediction capacity and empirical next-token prediction capacity and upper bounded them for general models. The rest of the paper will focus on lower bounding the next-token prediction capacity of the one-layer multi-head decoder-only transformer model, which we define in the next section. Before we move on though, we end the section with remarks on inexact interpolation, finite precision, and “VC dimension.”
Remark 14.
@@
Both Lemma 9 and Theorem 13 apply even for inexact interpolation. While Lemma 9 only uses that $F$, as defined in the proof, is not surjective, the proof actually shows that the image of $F$ has measure zero. If $\theta\mapsto h_{\theta}(\alpha)$ is additionally assumed to be $(\mathbb{R}_{\mathrm{an}},\exp)$-definable for all $\alpha\in\cup_{t=0}^{\infty}[\zeta]^{t}$, as it is in Theorem 13, then the image of $F$ is nowhere dense. Similarly, the image of $F$ defined in the proof of Theorem 13 is shown to be nowhere dense. So, for each choice of $k$ and $\zeta$, there exists $\epsilon>0$ such that the “$\epsilon$-approximate” next-token prediction capacity and “$\epsilon$-approximate” empirical next-token prediction capacity are both upper bounded by $O(k/\zeta)$.
Remark 15.
@@
Any next-token prediction capacity lower bound for a hypothesis class will apply even if the model is constrained to finite precision. To see this, suppose we have a hypothesis class $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ such that $\theta\mapsto h_{\theta}(\alpha)$ is continuously differentiable for all $\alpha\in\cup_{t=0}^{\infty}[\zeta]^{t}$. If we assume a maximum context length $T$ and restrict $\theta$ to a compact set, then the $\theta\mapsto h_{\theta}(\alpha)$ are uniformly Lipschitz continuous over $\alpha\in\cup_{t=0}^{\infty}[\zeta]^{t}$, so there exists $\delta>0$ such that any next-token prediction capacity lower bound also holds as an “$\epsilon$-approximate” next-token prediction capacity lower bound for the hypothesis class $\{h_{\theta}\mid\theta\in\mathbb{R}^{k}\}$ restricted to $\delta$-precision.
Remark 16.
It is interesting that the proof of Lemma 9 works for every distinct list $\alpha^{1},\ldots,\alpha^{n}$ while in the proof of Lemma 13 we have to construct a particular corpus $\beta^{1},\ldots,\beta^{N}$. In fact, we will show that there is another $\beta^{1},\ldots,\beta^{N}$ with $n$ unique contexts such that $\inf_{\theta\in\mathbb{R}^{k}}\mathcal{L}(\beta^{1},\ldots,\beta^{N},h_{\theta})=\mathcal{L}(\beta^{1},\ldots,\beta^{N},\hat{p})$. Note that this relates to the concept of Vapnik-Chervonenkis (VC) dimension [55]. If we define next-token prediction VC dimension as the largest $n$ such that there exists a distinct list $\alpha^{1},\ldots,\alpha^{n}$ such that for every list $y_{1},\ldots,y_{n}$ there exists $\theta$ such that $y_{i}=h_{\theta}(\alpha^{i})\leavevmode\nobreak\ \forall i\in[n]$ (a simple modification to Definition 7), then the same upper and lower bounds will hold as for next-token prediction capacity. On the other hand, if we define empirical next-token prediction VC dimension as the largest $n$ such that there exists $N$ and $\beta^{1},\ldots,\beta^{N}$ with $n$ unique contexts such that $\inf_{\theta\in\mathbb{R}^{k}}\mathcal{L}(\beta^{1},\ldots,\beta^{N},h_{\theta})=\mathcal{L}(\beta^{1},\ldots,\beta^{N},\hat{p})$ (a simple modification to Definition 10), then it will be infinite. Consider the model $h:\cup_{t=0}^{\infty}[\zeta]^{t}\to\mathrm{ri}\Delta^{\zeta-1}:\alpha\mapsto\mathbbm{1}/\zeta$. For any $T\in\mathbb{N}$, we can choose our text corpus to be $[\zeta]^{T}$. Then $n=(\zeta^{T+1}-1)/(\zeta-1)$ and $\mathcal{L}(\beta^{1},\ldots,\beta^{N},h)=\mathcal{L}(\beta^{1},\ldots,\beta^{N},\hat{p})$. Thus, if $h$ is in the hypothesis class, then the empirical next-token prediction VC dimension of the hypothesis class is infinity. Indeed, it is easy to see how to construct $h$ with the transformer model in the following section, proving that the transformer model has infinite empirical next-token prediction VC dimension.